OpenStudy (anonymous):

Determine whether the vectors u and v are parallel, orthogonal, or neither. u = <7, -4>, v = <-28, 16>

OpenStudy (anonymous):

@pizza12

OpenStudy (anonymous):

@wio @ganeshie8 @TuringTest @dan815

OpenStudy (anonymous):

I think parallel

OpenStudy (anonymous):

how'd you figure that out?

OpenStudy (anonymous):

I divided the numbers together

OpenStudy (dan815):

u.v=0 if orthogonal uXV=0 if parallel

OpenStudy (dan815):

u.v=|U||V|cos theta |u x v|= |U| |V| sin theta when they are orthogonal cos (90)=0 when they are parallel sin(0)=0

OpenStudy (anonymous):

well yeah I know that rule but I don't know the steps on how to calculate it unless its orthogonal...and I know it isn't orthogonal because it doesn't end up being 0 in the end

OpenStudy (dan815):

if u=k*v

OpenStudy (dan815):

then u know they are parallel

OpenStudy (anonymous):

take cross product \[u=-7i+4j,v=-28i+16j\] 

OpenStudy (dan815):

u = <7, -4>, v = <-28, 16> v=-4*(u)

OpenStudy (dan815):

therefore both u and v are vetors with same slope

OpenStudy (dan815):

just of different length

OpenStudy (dan815):

also given a vector u=<a,b> slope of u=b/a

OpenStudy (dan815):

slope of u=-4/7 slope of v=16/-28=4/-7=-4/7

OpenStudy (anonymous):

oh ok I think I forgot to write this part down in my notes:/ and I have one more question I don't understand I got wrong with the law of sines... Two triangles can be formed with the given information. Use the Law of Sines to solve the triangles. A = 56°, a = 16, b = 17 I keep getting -.5130898.... and I entered sin^-1 (16sin(56))/17)...what did I do wrong?

OpenStudy (anonymous):

nvm ill post this in another question:) thanks for all your help guys:)

OpenStudy (anonymous):